Your friend is a publisher. He tells you that if he raises the price of a book for 5 dollars (so the new price is 5 dollars higher than the old price), then his sales decrease for 100 copies (so the new sales is 100 copies fewer than old sales). He wants to know if he actually does this price change, what happens to his revenue (goes up, goes down or stay the same). What is your suggestion to him?
Answer #1

Suppose, my friend initially publishes X units of book and price charged is P per unit. Hence, total revenue=price*quantity = PX.

Now, let the new price be (P+5) per unit. As a result, quantity of books published becomes (X-100). New revenue = (P+5)(X-100) = PX + 5X - 100P - 500.

Change in revenue = PX+5X-100P-500-PX = 5X-100P-500

The change in revenue goes up , goes down or stay the same.

If 5X-100P-500 > 0, my friend will earn profit i.e, revenue goes up.

If 5X-100P-500 < 0, my friend will earn loss i.e, revenue goes down.

If 5X-100P-500 = 0, my friend will neither earn profit nor loss i,e, revenue remains unchanged.

To earn profit , my friend must produce at the level where 5X-100P-500 > 0 or, 5X-100P > 500 or, X - 20P > 100